# Metric Differential Privacy at the User-Level

### v1

@@
Abstract.
Metric differential privacy (DP) provides heterogeneous privacy guarantees based on a distance between the pair of inputs. It is a widely popular notion of privacy since it captures the natural privacy semantics for many applications (such as, for location data) and results in better utility than standard DP. However, prior work in metric DP has primarily focused on the item-level setting where every user only reports a single data item. A more realistic setting is that of user-level DP where each user contributes multiple items and privacy is then desired at the granularity of the user’s entire contribution. In this paper, we initiate the study of metric DP at the user-level. Specifically, we use the earth-mover’s distance ($d_{\textsf{EM}}$) as our metric to obtain a notion of privacy as it captures both the magnitude and spatial aspects of changes in a user’s data.
@@ Abstract.
User-level Differential Privacy, Earth-Mover’s Distance, Couplings
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacyccs: Theory of computation Design and analysis of algorithms
1. Introduction
@@
Intuitively, DP guarantee makes a pair of input data to be indistinguishable from each other. The standard DP guarantee requires all pairs of inputs to be indistinguishable thereby providing a uniform privacy guarantee to all pairs. This implies that every pair of input is considered equally sensitive. However, many practical applications call for a more tailored privacy semantics based on the heterogeneity of the data. In particular, input pairs that are closer or more similar to each other are considered to be more sensitive. For instance, for location data, revealing the exact city of residence is far more sensitive than revealing just the country. Metric DP ($d_{\mathcal{X}}$-DP; (Chatzikokolakis et al., 2013)) is a notion of DP that formally captures this heterogeneity in privacy semantics. Specifically, similarity is measured via a distance metric $d_{\mathcal{X}}$ and the privacy guarantee degrades linearly with the $d_{\mathcal{X}}$ distance between the pair of inputs. In addition to offering a more nuanced privacy definition, metric DP also improves utility compared to standard DP. This improvement stems from metric DP requiring only similar pairs of input to be indistinguishable, which results in a significantly lower noise than standard DP.
@@
Prior work in metric DP has primarily focused on the item-level setting where every user only reports a single data item (for e.g., a single record in a dataset). However, in many practical applications, a user contributes multiple items to a dataset. Privacy is then desired at the granularity of the user’s entire contribution. This has spurred a large body of work known as user-level DP (Amin et al., 2019; Bassily and Sun, 2023; Cummings et al., 2022; Acharya et al., 2023). However, all of this work considers only standard DP and is thus susceptible to the same limitations in utility as noted earlier. To this end, we initiate the study of metric DP at the user-level. While there have been some prior attempts at this, these work is limited to specific settings such as text data (Fernandes et al., 2019). To the best of our knowledge, this is the first work to give a general definition of metric DP at the user-level.
@@
The immediate task is to define a metric on the entire collection of a user’s data. Recall that metric DP caters to the privacy semantics that similar data is more sensitive. But the challenge here is that the similarity between two collections (sets) of data points has to be measured along two dimensions – $(1)$ the distance between the individual data items, and $(2)$ the fraction of the data items in the set that are different. In particular, note that in addition to small changes in the item-wise distances, changes in a smaller amount of the data also indicate more similarity and hence, correspond to more sensitive information (see below for concrete examples). This necessitates a measure that can express both of these quantities as a single metric. We tackle this challenge by using the earth-mover’s distance ($d_{\textsf{EM}}$(Givens and Shortt, 1984)) on the normalized representation of the user’s data. Informally, the $d_{\textsf{EM}}$ between two distributions is the minimum cost of transporting one distribution to another, where the cost is determined by the quantity of data items moved multiplied by the distance (measured via $d_{\mathcal{X}}$) over which they are moved. Our resulting privacy definition, denoted as $d_{\textsf{EM}}$-DP, yields the following privacy semantics. Under $d_{\textsf{EM}}$-DP, the strength of the privacy guarantee (indistinguishability) between two pairs of inputs $K,K^{\prime}$ (sets of data items) grows inversely with $\tau q$ if $K^{\prime}$ can be obtained by changing $\tau$ fraction of $K$ by an average distance of $q$ (Def. 3.1). $d_{\textsf{EM}}$ therefore takes into account both the structure of the distributions as well as the raw difference in their values. Consequently, the parameters $\tau$ and $q$ provide flexibility in interpretation and offer a nuanced privacy definition suitable for many practical applications. We illustrate this with the following examples:
@@
Textual Data. Consider a natural language dataset of user conversations where each user’s data is represented as a set of words. Typically, word embeddings $\phi$ map each word into a high-dimensional space, and word similarity is measured using a distance, such as the Euclidean distance, between $\phi(x_{1})$ and $\phi(x_{2})$. Now, the parameter $\tau$ corresponds to what fraction of the user’s conversation has changed in $K^{\prime}$ from $K$, while $q$ corresponds to the extent of the changes in the textual content. Thus, two conversations are harder to distinguish if $(1)$ there is only a fine-grained difference in their textual semantics  [^1] Such as transitioning from text about algebra to trigonometry versus changing it from ”math” to ”classical music”., and $(2)$ if it pertains to just a small fraction of the conversation (indicating a user rarely discussed the topic, which typically implies more sensitive information).
@@
Graph Data. Consider a graph $G=(V,E)$ in which connections in $E$ are private. Suppose there is additional public information in the form of a covariate $\phi:V\rightarrow\mathbb{R}^{d}$, which captures some auxiliary information about a user—for instance, the interests of a user. Here similarity between users is measured via covariate distance. The parameter $\tau$ corresponds to the fraction of a user’s connections which has changed in $K^{\prime}$ from $K$, and the parameter $q$ corresponds to the extent of the change in their interests. Thus, two graphs are harder to distinguish between if $(1)$ it is a fine-grained change to the interest222for instance, shifting from movies featuring Dwayne Johnson to Vin Diesel instead of from ”action” to “rom-com”, and $(2)$ if it pertains to only a few of the user’s connections. (say a small, private group of friends). This again captures natural privacy semantics as users are more likely to share common interests with their close friends than with a larger group, such as all workplace colleagues.
1.1. Details of Our Contributions
@@
We consider $n$ users who hold datasets $\{K_{i}\}_{i=1}^{n}$, each containing elements from a data domain $\mathcal{X}$ of size $k=|\mathcal{X}|$. Let $d_{\mathcal{X}}$ denote the distance metric defined over $\mathcal{X}$. WLOG, we consider $d_{\mathcal{X}}$ to be a normalized distance metric, i.e., all measures of distance are normalized to be at most $1$. Let $\tilde{K}_{i}$ denote the normalized version of the dataset $K_{i}$. $d_{\textsf{EM}}$ between any pair of datasets $\{K_{i},K_{i}^{\prime}\}$ can be defined by first normalizing them to $\{\tilde{K}_{i},\tilde{K}_{i}^{\prime}\}$, and then using $d_{\mathcal{X}}$ to measure the minimum cost of transporting $\tilde{K}_{i}$ to $\tilde{K}_{i}^{\prime}$. The global dataset is given by $K_{G}=K_{1}\cup\cdots\cup K_{n}$, and there is an aggregator who wants to privately compute a query $F(K)$. In the central model, the aggregator already holds $K_{i}$ from each user, and applies a private mechanism $\mathcal{M}(K)$ to obtain a private estimate for $F$. In the local model, the users do not trust the aggregator, and communicate private messages $\{m_{i}=\mathcal{M}_{i}(K_{i})\}$ to the aggregator. The aggregator then post-processes these messages $\mathcal{F}(m_{1},\ldots,m_{n})$ to output a private estimate of $F$. For simplicity, in this work we assume the mechanisms $\mathcal{M}_{i}$ to be non-interactive.
@@
We also make a distinction between bounded and unbounded data. Note that boundedness here refers to the size of each user’s dataset and not the number of the users – throughout the paper, we assume that the number of users, $n$, is fixed and publicly known. In our specific context, bounded data corresponds to the case where the size of each user’s dataset is publicly known, and the mechanism $\mathcal{M}$ only needs to preserve privacy between datasets of the same size. Furthermore, in the central model, each user’s dataset has the same public size. The benefit of this simplification is that algorithm analysis is easier. Such a bounded data setting has been considered in many previous works (Li et al., 2016). We also consider the general unbounded data setting where each user can have datasets of varying sizes, with the size being private as well.
@@
For each model and type of boundedness, we summarize how one would apply $d_{\textsf{EM}}$-DP, along with the resulting semantics, in Table 1. We also include a corresponding notion of the standard user-level DP (Liu et al., 2023) (provides a uniform privacy guarantee to all pairs of datasets) and serves as our baseline. In what follows, we elaborate on our main contributions.
@@
Local (applies to each $\mathcal{M}_{i}$) User Unbounded $(\varepsilon,\delta)$-user-level DP (Def. 2.1) Two input datasets $K,K^{\prime}\in\mathcal{X}^{*}$ are indistinguishable with parameters $(\varepsilon,\delta)$ Recently proposed in (Acharya et al., 2023). Acts our baseline for the local model.
@@
Item N/A $(\alpha,\delta)$-$d_{\mathcal{X}}$-DP (Def. 2.3) Two input items $x,x^{\prime}\in\mathcal{X}$ is protected with parameters $(\alpha d_{\mathcal{X}}(x,x^{\prime}),\delta)$ Proposed in (Chatzikokolakis et al., 2015)
Central (applies to $\mathcal{M}$) User Unbounded $(\varepsilon,\delta)$-user-level DP (Def. 2.2) Let $K_{G}=K_{1}\cup\cdots K_{n}$ where $K_{i}\in\mathcal{X}^{*}$. Two input global datasets $K_{G},K_{G}^{\prime}$ s.t. they differ only on the dataset of a single user $\{K_{i},K_{i}^{\prime}\},i\in[n]$ are indistinguishable with parameters $(\varepsilon,\delta)$ Studied widely (Bassily and Sun, 2023; Liu et al., 2020, 2023). Acts our baseline for the central model.
User Bounded $(\alpha,\delta)$-bounded $d_{\textsf{EM}}$-DP (Def. 3.1) Two input global datasets $K_{G},K_{G}^{\prime}$ s.t. they differ only on $\{K_{i},K_{i}^{\prime}\}\in\mathcal{X}^{m}\times\mathcal{X}^{m}$ are indistinguishable with parameters $(\alpha d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime}),\delta)$ Each $K_{i}$ has size $m$ which is public.
User Unbounded $(\varepsilon,\delta,r)$-discrete $d_{\textsf{EM}}$-DP (Def. 3.2) Two input global datasets $K_{G},K_{G}^{\prime}$ s.t. they differ only on $\{K_{i},K_{i}^{\prime}\}$ and $d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})\leq r$ are indistinguishable with parameters $(\varepsilon,\delta)$. Using group privacy, we can show the following parameters $(\varepsilon\lceil\frac{d}{r}\rceil,\delta\exp(\varepsilon\lceil\frac{d}{r}\rceil))$ where $d=d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})$. Implies user-level DP when $r\geq 1$ since $d_{\textsf{EM}}(\cdot,\cdot)\leq 1$.
Table 1. Summary of privacy definitions for this paper. The number of users, $n$, is fixed and publicly known for all the definitions.
1.1.1. Mechanism Design
We provide novel mechanisms for answering two types of queries for $d_{\textsf{EM}}$-DP.

### v2

@@
Abstract
Metric differential privacy (DP) provides heterogeneous privacy guarantees based on a distance between the pair of inputs. It is a widely popular notion of privacy since it captures the natural privacy semantics for many applications (such as, for location data) and results in better utility than standard DP. However, prior work in metric DP has primarily focused on the item-level setting where every user only reports a single data item. A more realistic setting is that of user-level DP where each user contributes multiple items and privacy is then desired at the granularity of the user’s entire contribution. In this paper, we initiate the study of one natural definition of metric DP at the user-level. Specifically, we use the earth-mover’s distance ($d_{\textsf{EM}}$) as our metric to obtain a notion of privacy as it captures both the magnitude and spatial aspects of changes in a user’s data.
@@ Abstract
1 Introduction
@@
Intuitively, DP guarantee makes a pair of input data to be indistinguishable from each other. The standard DP guarantee requires all pairs of inputs to be indistinguishable thereby providing a uniform privacy guarantee to all pairs. This implies that every pair of input is considered equally sensitive. However, many practical applications call for a more tailored privacy semantics based on the heterogeneity of the data. In particular, input pairs that are closer or more similar to each other are considered to be more sensitive. For instance, for location data, revealing the exact city of residence is far more sensitive than revealing just the country. Metric DP ($d_{\mathcal{X}}$-DP; Chatzikokolakis et al. (2013)) is a notion of DP that formally captures this heterogeneity in privacy semantics. Specifically, similarity is measured via a distance metric $d_{\mathcal{X}}$ and the privacy guarantee degrades linearly with the $d_{\mathcal{X}}$ distance between the pair of inputs. In addition to offering a more nuanced privacy definition, metric DP also improves utility compared to standard DP. This improvement stems from metric DP requiring only similar pairs of input to be indistinguishable, which results in a significantly lower noise than standard DP.
@@
Prior work in metric DP has primarily focused on the item-level setting where every user only reports a single data item (for e.g., a single record in a dataset). However, in many practical applications, a user contributes multiple items to a dataset. Privacy is then desired at the granularity of the user’s entire contribution. This has spurred a large body of work known as user-level DP (Amin et al., 2019; Bassily and Sun, 2023; Cummings et al., 2022; Acharya et al., 2023). However, all of this work considers only standard DP and is thus susceptible to the same limitations in utility as noted earlier. To this end, we initiate the study of one natural definition of metric DP at the user-level. In particular, we look into a specific distance metric called earth-mover’s distance ($d_{\textsf{EM}}$ivens and Shortt (1984)). $d_{\textsf{EM}}$ measures the similarity of two distributions and is quantified by the cost of transforming one distribution to another where the cost function can be defined by any metric over the underlying data. Thus, $d_{\textsf{EM}}$ provides a general and naturally interpretable way of measuring similarity, suitable for capturing the privacy semantics of various real-world scenarios. While there have been some prior attempts at this, these works are limited to specific settings, such as text data Fernandes et al. (2019). To the best of our knowledge, this is the first work to propose a definition of metric DP at the user-level for a general setting via $d_{\textsf{EM}}$.
@@
The immediate question when applying metric DP at the user level is how to define a metric on the entire collection of a user’s data. We argue that $d_{\textsf{EM}}$ is particularly well-suited for this task. Recall that metric DP caters to the privacy semantics that similar data is more sensitive. But the challenge here is that the similarity between two collections (sets) of data points has to be measured along two dimensions – $(1)$ the distance between the individual data items, and $(2)$ the fraction of the data items in the set that are different. In particular, note that in addition to small changes in the item-wise distances, changes in a smaller amount of the data also indicate more similarity and hence, correspond to more sensitive information (see below for concrete examples). This necessitates a measure that can express both of these quantities as a single metric, for which $d_{\textsf{EM}}$ is a natural choice. Informally, the $d_{\textsf{EM}}$ between two distributions is the minimum cost of transporting one distribution to another, where the cost is determined by the quantity of data items moved multiplied by the distance (measured via $d_{\mathcal{X}}$) over which they are moved. Our resulting privacy definition, denoted as $d_{\textsf{EM}}$-DP, yields the following privacy semantics. Under $d_{\textsf{EM}}$-DP, the strength of the privacy guarantee (indistinguishability) between two pairs of inputs $K,K^{\prime}$ (sets of data items) grows inversely with $\tau q$ if $K^{\prime}$ can be obtained by changing $\tau$ fraction of $K$ by an average distance of $q$ (Def. 3.1). $d_{\textsf{EM}}$ therefore takes into account both the structure of the distributions as well as the raw difference in their values. Consequently, the parameters $\tau$ and $q$ provide flexibility in interpretation and offer a nuanced privacy definition suitable for many practical applications. We illustrate this as follows:
@@
Textual Data. Consider a natural language dataset of user conversations where each user’s data is represented as a set of words. Typically, word embeddings $\phi$ map each word into a high-dimensional space, and word similarity is measured using a distance, such as the Euclidean distance, between $\phi(x_{1})$ and $\phi(x_{2})$. Now, the parameter $\tau$ corresponds to what fraction of the user’s conversation has changed in $K^{\prime}$ from $K$, while $q$ corresponds to the extent of the changes in the textual content. Thus, two conversations are harder to distinguish if $(1)$ there is only a fine-grained difference in their textual semantics [^1] Such as transitioning from text about algebra to trigonometry versus changing it from “math” to “classical music”., and $(2)$ if it pertains to just a small fraction of the conversation (indicating a user rarely discussed the topic, which typically implies more sensitive information).
@@
Graph Data. Consider a graph $G=(V,E)$ in which connections in $E$ are private. Suppose there is additional public information in the form of a covariate $\phi:V\rightarrow\mathbb{R}^{d}$, which captures some auxiliary information about a user—for instance, the interests of a user. Here similarity between users is measured via covariate distance. The parameter $\tau$ corresponds to the fraction of a user’s connections which has changed in $K^{\prime}$ from $K$, and the parameter $q$ corresponds to the extent of the change in their interests. Thus, two graphs are harder to distinguish between if $(1)$ it is a fine-grained change to the interest222for instance, shifting from movies featuring Dwayne Johnson to Vin Diesel instead of from “action” to “rom-com”, and $(2)$ if it pertains to only a few of the user’s connections (say a small, private group of friends). This again captures natural privacy semantics as users are more likely to share common interests with their close friends than with a larger group, such as all workplace colleagues.
In a nutshell, $d_{\textsf{EM}}$-DP offers a more fine-grained privacy definition compared to standard DP, that captures real-world privacy semantics more effectively while providing better utility. For example, consider the following two instances of $K^{\prime}$ in the aforementioned example of location data – one where the data for an entire month is different and another where only a single day’s data differs. Standard DP treats both cases as equally sensitive (i.e., offers the same privacy guarantee for both cases), necessitating a larger noise addition even in the latter case, which results in reduced utility. In contrast, $d_{\textsf{EM}}$-DP offers a stronger privacy guarantee for the latter, thereby resulting in a better privacy-utility trade-off. $d_{\textsf{EM}}$-DP is in fact a natural relaxation of standard DP. We provide more details on the interpretation of $d_{\textsf{EM}}$-DP relative to standard DP in Sec. 8. A full version of this paper appears in Imola et al. (2024).
1.1 Details of Our Contributions
@@
We consider $n$ users who hold datasets $\{K_{i}\}_{i=1}^{n}$, each containing elements from a data domain $\mathcal{X}$ of size $k=|\mathcal{X}|$. Let $d_{\mathcal{X}}$ denote a distance metric defined over $\mathcal{X}$. WLOG, we consider $d_{\mathcal{X}}$ to be a normalized distance metric, i.e., all measures of distance are normalized to be at most $1$. Let $\tilde{K}_{i}$ denote the normalized version of the dataset $K_{i}$. $d_{\textsf{EM}}$ between any pair of datasets $\{K_{i},K_{i}^{\prime}\}$ can be defined by first normalizing them to $\{\tilde{K}_{i},\tilde{K}_{i}^{\prime}\}$, and then using $d_{\mathcal{X}}$ to measure the minimum cost of transporting $\tilde{K}_{i}$ to $\tilde{K}_{i}^{\prime}$. The global dataset is given by $K_{G}=K_{1}\cup\cdots\cup K_{n}$, and there is an aggregator who wants to privately compute a query $V(K)$. In the central model, the aggregator already holds $K_{i}$ from each user, and applies a private mechanism $\mathcal{M}(K_{G})$ to obtain a private estimate for $V$. In the local model, the users do not trust the aggregator, and communicate private messages $\{m_{i}=\mathcal{M}_{i}(K_{i})\}$ to the aggregator. The aggregator then post-processes these messages $\mathcal{V}(m_{1},\ldots,m_{n})$ to output a private estimate of $V$. For simplicity, in this work we assume the mechanisms $\mathcal{M}_{i}$ to be non-interactive.
@@
We also make a distinction between bounded and unbounded data. Note that boundedness here refers to the size of each user’s dataset and not the number of the users – throughout the paper, we assume that the number of users, $n$, is fixed and publicly known. In our specific context, bounded data corresponds to the case where the size of each user’s dataset is publicly known, and the mechanism $\mathcal{M}$ only needs to preserve privacy between datasets of the same size. Furthermore, in the central model, each user’s dataset has the same public size. The benefit of this simplification is that algorithm analysis is easier. Such a bounded data setting has been considered in many previous works (see  Li et al. (2016)). We also consider the general unbounded data setting where each user can have datasets of varying sizes, with the size being private as well.
@@
For each model and type of boundedness, we summarize how one would apply $d_{\textsf{EM}}$-DP, along with the resulting semantics, in Table 1. We also include a corresponding notion of the standard user-level DP Liu et al. (2023) (provides a uniform privacy guarantee to all pairs of datasets) which serves as our baseline. In what follows, we elaborate on our main contributions.
@@
Local (applies to each $\mathcal{M}_{i}$) User Unbounded $(\varepsilon,\delta)$-user-level DP (Def. 2.1) Two input datasets $K,K^{\prime}\in\mathcal{X}^{*}$ are indistinguishable with parameters $(\varepsilon,\delta)$ Recently proposed in Acharya et al. (2023). Acts our baseline for the local model.
@@
User Unbounded $(\varepsilon,\delta,r)$-discrete $d_{\textsf{EM}}$-DP (Def. 5.1) Two input datasets $K,K^{\prime}\in\mathcal{X}^{*}$ such that $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\leq r)$ are indistinguishable with parameters $(\varepsilon,\delta)$ Using group privacy, can show $(\varepsilon\lceil\frac{d}{r}\rceil,\delta\exp(\varepsilon\lceil\frac{d}{r}\rceil)$ for two $K,K^{\prime}$ s.t. $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\leq d$.
Item N/A $(\alpha,\delta)$-$d_{\mathcal{X}}$-DP (Def. 2.3) Two input items $x,x^{\prime}\in\mathcal{X}$ is protected with parameters $(\alpha d_{\mathcal{X}}(x,x^{\prime}),\delta)$ Proposed in Chatzikokolakis et al. (2015).
Central (applies to $\mathcal{M}$) User Unbounded $(\varepsilon,\delta)$-user-level DP (Def. 2.2) Let $K_{G}=K_{1}\cup\cdots K_{n}$ where $K_{i}\in\mathcal{X}^{*}$. Two input global datasets $K_{G},K_{G}^{\prime}$ s.t. they differ only on the dataset of a single user $\{K_{i},K_{i}^{\prime}\},i\in[n]$ are indistinguishable with parameters $(\varepsilon,\delta)$ Studied widely Bassily and Sun (2023); Liu et al. (2020, 2023). Acts our baseline for the central model.
User Bounded $(\alpha,\delta)$-bounded $d_{\textsf{EM}}$-DP (Def. 3.2) Two input global datasets $K_{G},K_{G}^{\prime}$ s.t. they differ only on $\{K_{i},K_{i}^{\prime}\}\in\mathcal{X}^{m}\times\mathcal{X}^{m}$ are indistinguishable with parameters $(\alpha d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime}),\delta)$ Each $K_{i}$ has size $m$ which is public.
User Unbounded $(\varepsilon,\delta,r)$-discrete $d_{\textsf{EM}}$-DP (Def. 5.2) Two input global datasets $K_{G},K_{G}^{\prime}$ s.t. they differ only on $\{K_{i},K_{i}^{\prime}\}$ and $d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})\leq r$ are indistinguishable with parameters $(\varepsilon,\delta)$. Using group privacy, can show parameters $(\varepsilon\lceil\frac{d}{r}\rceil,\delta\exp(\varepsilon\lceil\frac{d}{r}\rceil))$ for any $K_{i},K_{i}^{\prime}$ s.t. $d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})\leq d$. Implies user-level DP when $r\geq 1$.
Table 1: Summary of privacy definitions for this paper. The number of users, $n$, is fixed and publicly known for all the definitions.
1.1.1 Mechanism Design
We provide novel mechanisms for answering two types of queries for $d_{\textsf{EM}}$-DP. In this section, we let $K$ denote a general dataset of interest, which in the local model would be set to $K_{i}$ and in the global model would be set to $K_{G}$.